Arthur can edit 15 essays in 10 h. Debra needs 15 h to do the same job. If they work together, how long will it take them to edit the essays?
Answer by stanbon(75887)

Arthur can edit 15 essays in 10 h. Debra needs 15 h to do the same job. If they work together, how long will it take them to edit the essays?
-----------------
Arthur DATA:
time = 10 hr/job ; rate = 1/10 job/hr
-------
Debra DATA:
time = 15 hr/job ; rate = 1/15 job/hr
--------
Together DATA:
time = x hr/job ; rate = 1/x job/hr
==================
Equation:
rate + rate = together rate
1/10 + 1/15 = 1/x
---
Multiply thru by 30x to get:
3x + 2x = 30
5x = 30
x = 6 hrs (amt. of time needed for them to do the job together)
========================
Cheers,
Stan H.
===============